OpenStudy (anonymous):

Find the sum. 5x/7x-y + 4/y-7x plz help

OpenStudy (misty1212):

lol that is not right!!

OpenStudy (misty1212):

is it \[\huge \frac{5x}{7x-y}+\frac{4}{y-7x}\]?

OpenStudy (sweetburger):

get teh same denominator

OpenStudy (anonymous):

Get to same denominator like this: \[(\frac{ 5x }{ 7x-y }\times \frac{ y-7x }{ y-7x }) + (\frac{ 4 }{ y-7x } \times \frac{ 7x-y }{ 7x-y }) = \frac{ 5xy-35x }{ 14xy-49x ^{2} -y ^{2} }+ \frac{ 28x-4y }{ 14xy-49x ^{2}-y ^{2} }\]